OpenStudy (anonymous):

The height in meters of a projectile can be modeled by h = −4.9t2 + vt + s where t is the time (in seconds) the object has been in the air, v is the initial velocity (in meters per seconds) and s is the initial height (in meters). A soccer ball is kicked upward from the ground and flies through the air with an initial vertical velocity of 4.9 meters per second. After how many seconds does it land approximately?

OpenStudy (anonymous):

\[-4.9t^2+4.9t=0\] solve for \(t\)

OpenStudy (anonymous):

factor as \[4.9t(t-1)=0\] then it is pretty clear what \(t\) has to be right?

OpenStudy (anonymous):

if not, let me know

OpenStudy (anonymous):

ya, sorry, i'm still confused.....

OpenStudy (anonymous):

if \(t-1=0\) then \(t=1\)

OpenStudy (anonymous):

what might be confusing is what \(v\) is in this example \(v=4.9\)

OpenStudy (anonymous):

yeah, that was throwing me off, thank you

OpenStudy (anonymous):

i thought it might be yw